OpenStudy (anonymous):

Use intercepts to graph each line by each equation. 1) 2x-4y=16 2) -3y+6x=(-18) 3) y=-3x+3

OpenStudy (anonymous):

Well you can begin by recalling that you can obtain the x intercept, by setting y to be zero and solving for x. Likewise you can get the y intercept by setting x to be zero and solving for y

OpenStudy (anonymous):

Ah, thank you.

OpenStudy (amistre64):

intercepts are fun if; I do it like this: given a linear equation: ax + by = c, we can divide out the c \[\frac{ax}{c}+\frac{by}{c}=\frac{c}{c}\]then we can divide off the numerators as well:\[\frac{x}{c/a}+\frac{y}{c/b}=1\] and your intercepts are just the denominators now

OpenStudy (amistre64):

2x-4y=16 2x/16 -4y/16 = 16/16 x/16/2 + y/16/-4 = 1 x/8 + y/-4 = 1 you xint = 8 and your yint = -4

OpenStudy (amistre64):

-3y + 6x = (-18) -3y/(-18) + 6x/(-18) = (-18)/(-18) y/-18/-3 + x/-18/6 = 1 y/6 + x/-3 = 1 xint=-3, yint=6